OpenStudy (naveenbhatia1312):

Find an equation for the nth term of the arithmetic sequence. a19 = -92, a20 = 6 an = -1856 + 98(n - 1) an = -1856 - 98(n - 1) an = -1856 - 98(n + 1) an = -1856 + 98(n + 1)

Directrix (directrix):

an=a₁+(n-1)*d Formula for nth term --------------- a19 = -92 -92 = a₁ + 18*d ------- a20 = 6 6 = a₁ + 19*d ----- Solve these two equations for d and a₁ -92 = a₁ + 18*d 6 = a₁ + 19*d --------------

OpenStudy (naveenbhatia1312):

im confused

OpenStudy (naveenbhatia1312):

@Directrix

Directrix (directrix):

Add the second equation from the first (the ones on the right side) so that you can solve for d which is the common difference of the terms in the sequence. -92 = a₁ + 18*d ---> -92 = a₁ + 18*d 6 = a₁ + 19*d -- (-1)--> -6 = - a₁ -19*d =============